OpenStudy (anonymous):

A line has a slope of and a run = 50. What is the rise? 25 50 100

Nnesha (nnesha):

slope `of` -what-- ? slope = rise over run

Nnesha (nnesha):

retype ur question

OpenStudy (anonymous):

thats the question..

OpenStudy (anonymous):

oh its the slope os 1/2

Nnesha (nnesha):

slope = rise over run given is run we need to find `rise` so \[\huge\rm \color{reD}{slope} =\frac{ rise }{\color{blue}{ run}}\] \[\huge\rm \color{reD}{\frac{1}{2}} =\frac{ rise }{\color{blue}{50}}\] now solve for rise